# Can Knowledge Graphs Make Large Language Models More Trustworthy? An Empirical Study Over Open-ended Question Answering

### v3

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have led to promising improvements in enhancing the reasoning accuracy of Large Language Models (LLMs). However, current benchmarks focus mainly on closed-ended tasks, leaving a gap in the assessment of more complex real-world scenarios. This gap has also obscured the evaluation of KGs’ potential to mitigate the problem of hallucination in LLMs. To fill the gap, we introduce OKGQA, a new benchmark specifically designed to assess LLMs enhanced with KGs under open-ended, real-world question answering scenarios. OKGQA is designed to closely reflect the complexities of practical applications using questions from different types, and incorporates specific metrics to measure both hallucination ratio and the enhancement in reasoning capabilities. To consider the scenario in which KGs may have varying levels of mistakes, we propose another benchmark variant OKGQA-P to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. OKGQA aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe that this study can facilitate a more complete performance comparison and encourage continuous improvement in integrating KGs with LLMs to reduce hallucination.
@@ 1 Introduction
Contemporary LLMs are prone to producing hallucinations due to gaps in their knowledge Gekhman et al. (2024); Lee et al. (2023). These inaccuracies commonly stem from misinformation, biases, or errors in the training data, and lead to responses that seem plausible but may be irrelevant or incorrect Weng (2024). This issue is particularly concerning in high-stakes contexts such as healthcare He et al. (2023) and science Taylor et al. (2022) [^1] Code and data are released at https://anonymous.4open.science/r/OKGQA-CBB0.
@@
Figure 1: Comparison between Open-ended Question and Close-ended Question over Knowledge Graph.
To address this limitation, researchers have turned to leveraging external knowledge graphs (KGs) as a complementary Pan et al. (2024); Luo et al. (2023a); Hu et al. (2023); Sui et al. (2024). KGs offer structured and explicit factual information—often domain-specific—and allow each piece of data to be traced back to its source Zheng et al. (2023); Agrawal et al. (2023). This traceability not only enables verification of the model’s reasoning but also brings transparency to the decision-making process, making KGs a promising method for enhancing the reliability of LLM reasoning. Find more details of related works in §C.
@@
However, current benchmarks for testing these LLM+KG models are predominantly closed-ended Jin et al. (2020); Puerto et al. (2023), restricting the model’s output to a limited set of entities, relations Talmor et al. (2019); Mihaylov et al. (2018) or logical forms Yih et al. (2016); Talmor and Berant (2018). While these benchmarks are useful to measure retrieval and basic reasoning, they do not adequately capture whether a model is hallucinating. In closed-ended settings, errors can stem from incorrect retrieval or from fabricating (hallucinating) answers, yet conventional metrics (e.g., accuracy or precision) cannot distinguish between these two issues. This becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al. (2022).
@@
In contrast, our work focuses on open-ended KGQA, where LLMs are prompted to generate more elaborate answers, include reasoning paths and supporting facts derived from the KG (as shown in Figure 1). This broader output space offers two key advantages: First, it enables direct assessments of hallucination with metrics like FActScore Min et al. (2023) or SAFE Wei et al. (2024)), which decompose longer responses into atomic statements for factual consistency checks with external knowledge sources like Wikipedia. Second, it increases the likelihood of exposing factual errors which helps to assess the phenomenon indicated in Qiu et al. (2024): where longer, more complex responses provide more opportunities for errors to occur. By adopting this open-ended approach, we aim to (1) explore whether KGs can make LLMs more trustworthy in the open-ended setting, and (2) conduct a comparative analysis to shed light on methods design and direction for leveraging KGs to reduce LLMs’ hallucination.
@@
To achieve this, we introduce a new benchmark, Open-ended Knowledge-Graphs Question Answering (OKGQA), specifically designed to assess LLMs enhanced with KGs in an open-ended QA setting. OKGQA is designed to closely reflect the complexities of practical applications with diverse questions types as mentioned in Table 1, ensuring that all the queries cannot be answered simply by retrieving isolated KG facts. To consider the scenarios for potentially contaminated or imperfect KGs (i.e., attributes may be mislabeled, relations may not exist, etc.), we also propose a variant OKGQA-P (§2.2) to assess model performance under conditions where KGs’ semantics and structure are deliberately perturbed and contaminated. In both settings, we measure the degree of hallucination and the overall quality of the responses (see §4.1 for details).
@@
Based on our experiments, we find that (1) integrating KG information generally mitigates factual errors, especially for queries requiring deep reasoning; (2) directly performing reasoning in the LLM itself (e.g., internal reasoning strategies like Chain-of-thought Kim et al. (2023) and Self-Consistency Wang et al. (2022)) may cause bias and hallucination; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
@@ 2 OKGQA: An Open-ended Knowledge Graph Question-Answering Benchmark
OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. As illustrated in Figure 1, given a complex query and its corresponding subgraph in a KG, the system must be capable of understanding the relationships within the data and performing human-like reasoning over the KG content to compose a paragraph-long answer. In the following section, we first describe our dataset construction, including query generation via LLM templates and KG subgraph extraction with PPR pruning. We then introduce OKGQA-P, a benchmark variant designed to evaluate model robustness under KG perturbations, detailing our perturbation methods and the metrics used to assess semantic and structural deviations. Due to the page limitation, we also include some extension of our benchmark in Appendix §B, including multilingual setup, and more analysis.
@@ 2.1 Dataset Construction
We utilize a template-based approach to generate a diverse range of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Details regarding specific templates and example queries can be found in Table 1, while the corresponding prompts are provided in the Appendix D. To ensure that the generated queries represent real-world scenarios and complexities, we employ an iterative optimization approach that utilizes both automated and human evaluation to refine the query generation process (the details are given in Appendix A.1). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates undergo automated evaluation using an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, with higher scores indicating better performance across multiple metrics. Subsequently, human evaluators assess the same queries, producing corresponding normalized scores $s_{\text{human}}$ within the same range. To refine the dataset, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$. This optimization process ensures alignment between automated and human quality assessments. The queries are also categorized by complexity with detailed statistics in Figure 2.
@@
To reduce the size of KGs while retaining relevant information, we follow previous work Yih et al. (2016); Talmor and Berant (2018) by sampling subgraphs from DBpedia (Noted that all queries in OKGQA can be answered using DBpedia). We extract all triples contained within the $K$-hop neighbors from the entities mentioned in the query. We set $K=2$ to balance graph size and computational feasibility. As increasing beyond $2$-hop subgraphs generally leads to exponential growth in edges and nodes Jin et al. (2020), which increase excessive noise and complicating information retrieval [^2] This choice is also informed by common practices in other benchmarks, such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of the 2-hop subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al. (2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.2). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
@@ 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) for capturing semantics (ATS) and structural (SC2D, SD2) similarity between KGs. Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each of the three metrics, higher value indicates higher similarity. The detailed description can be found in Appendix A.5, with visualization in Figure 5.
@@ 3.1 Formalization
We formalize the KG-augmented framework for reducing hallucination as follows. Given a user query $q$, a pretrained language model generates a paragraph-like answer $a$ by modeling the conditional probability $p(a|q)$. To explore whether KGs help reduce hallucinations of LLMs, we introduce retrieved knowledge $\mathcal{Z}$ from the KG and define:
@@ 3.2 Graph-guided retrieval (G-retrieval)
  • Sub-graph retrieval: building on He et al. (2024a), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
@@ 4 Experiments
In this section, we first introduce the evaluation metrics, and then focus on two main research questions: RQ1: Can KGs reduce hallucination in LLMs? and RQ2: How are KG-Aware methods affected by noise/perturbations in KGs?
@@ 4.1 Evaluation Metrics & Setup
In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 2024) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023). In specific, we use G-Eval (Liu et al., 2023) framework for the evaluation and provide relevant Wikipedia pages of each query as context to enhance G-Eval’s robustness and stability. The four metrics are defined as follows: Context Relevance: measures how well the generated response aligns with the provided context. Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. Correctness: measures the clarity and specificity of the generated answer to the question. Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The detailed prompt can be found in Appendix D.
We use gpt-4o-mini (from November 2024 to January 2025) as LLM backbone for all the evaluation metrics. As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the models Gu et al. (2024); Li et al. (2024), we conduct additional validation in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLM in the LLM-as-evaluator framework has little impact on the overall evaluation and demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
For other testing LLM backbones mentioned in this section, we consider a range of widely used LLMs of different scales, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). Considering the trade-off between cost and performance, we use text-embedding-3-small model from OpenAI (from November 2024 to January 2025) as embedding model for G-retrieval process.
@@ 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency, as shown in Var-1 in Table 2. It shows that these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish. This shows that relying solely on internal reasoning is inadequate for mitigating hallucinations, highlighting the necessity for external knowledge to address this issue effectively.
@@ 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We benchmark different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings.
Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%. In summary, we find that effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. Considering practical scenarios, platforms like Wikidata are less likely to have perturbations as severe as 50% due to their ongoing updates and community-based quality control. This ensures the relevance and applicability of our findings in real-world settings.

### v4

@@ Abstract
Recent works integrating Knowledge Graphs (KGs) have shown promising improvements in enhancing the reasoning capabilities of Large Language Models (LLMs). However, existing benchmarks primarily focus on closed-ended tasks, leaving a gap in evaluating performance on more complex, real-world scenarios. This limitation also hinders a thorough assessment of KGs’ potential to reduce hallucinations in LLMs. To address this, we introduce OKGQA [^1] Code and data are released at https://github.com/Y-Sui/OKGQA, a new benchmark specifically designed to evaluate LLMs augmented with KGs in open-ended, real-world question answering settings. OKGQA reflects practical complexities through diverse question types and incorporates metrics to quantify both hallucination rates and reasoning improvements in LLM+KG models. To consider the scenarios in which KGs may contain varying levels of errors, we propose a benchmark variant, OKGQA-P, to assess model performance when the semantics and structure of KGs are deliberately perturbed and contaminated. In this paper, we aims to (1) explore whether KGs can make LLMs more trustworthy in an open-ended setting, and (2) conduct a comparative analysis to shed light on method design. We believe this study can facilitate a more complete performance comparison and encourages continuous improvement in integrating KGs with LLMs to mitigate hallucination, and make LLMs more trustworthy.
@@ 1 Introduction
Contemporary large language models (LLMs) are prone to producing hallucinations—plausible-sounding but incorrect or irrelevant outputs—due to gaps and inconsistencies in their training data Gekhman et al. (2024); Lee et al. (2022). These inaccuracies often arise from misinformation, biases, or errors embedded in the data Weng (2024); Chen et al. (2025), posing significant risks in high-stakes domains such as healthcare He et al. (2023b) and scientific research Taylor et al. (2022).
@@
Figure 1: Comparison between open-ended and closed-ended questions over knowledge graphs.
To mitigate these limitations, recent research has explored augmenting LLMs with external knowledge graphs (KGs) Pan et al. (2024); Luo et al. (2023a); Hu et al. (2023); Sui et al. (2024). KGs provide structured, explicit, and often domain-specific factual information, with each fact traceable to its original source Zheng et al. (2023); Agrawal et al. (2023). These properties not only enable verification of the model’s reasoning but also bring transparency to the decision-making process, making KGs a promising avenue for improving the reliability and trustworthiness of LLM outputs. A detailed review of related work is provided in Appendix §C.
@@
However, existing benchmarks for evaluating these LLM+KG models predominantly focus on closed-ended tasks Jin et al. (2020); Puerto et al. (2021), where model outputs are restricted to a fixed set of entities, relations Talmor et al. (2019); Mihaylov et al. (2018), or logical forms Yih et al. (2016); Talmor and Berant (2018). While these benchmarks are useful to measure retrieval accuracy and basic reasoning, such benchmarks fall short in detecting whether a model is hallucinating. In closed-ended settings, errors can stem from incorrect retrieval or fabricated (hallucinated) answers, but conventional metrics like accuracy or precision cannot distinguish between these two failure modes. This limitation becomes problematic for more complex, real-world applications that demand nuanced answers Kantharaj et al. (2022).
@@
In contrast, our work focuses on open-ended knowledge graph question answering (KGQA), where LLMs are aimed to generate detailed answers that include explicit reasoning paths and supporting facts derived from the KG, as illustrated in Figure 1. This expanded output space offers two key advantages. First, it enables direct measurement of hallucination using metrics such as FActScore Min et al. (2023) and SAFE Wei et al. (2024), which decompose complex responses into atomic statements and verify their factual consistency against external knowledge sources like Wikipedia. Second, longer and more complex answers increase the likelihood of exposing factual errors, aligning with observations from Qiu et al. (2024) that hallucinations tend to accumulate in extended responses. By adopting this open-ended paradigm, we aim to (1) explore whether KGs can enhance the trustworthiness of LLMs in realistic, open-ended scenarios, and (2) provide a comparative analysis to guide the design of methods that leverage KGs to reduce hallucination.
@@
To this end, we introduce Open-ended Knowledge-Graphs Question Answering (OKGQA), a new benchmark tailored to evaluate LLMs augmented with KGs in open-ended QA settings. OKGQA reflects the complexities of practical applications by incorporating diverse question types (see Table 1), ensuring that queries cannot be answered by simply retrieving isolated KG facts. To simulate real-world conditions where KGs may be imperfect or contaminated—for example, with mislabeled attributes or spurious relations—we propose a benchmark variant, OKGQA-P (§2.2), which evaluates model robustness when KG semantics and structure are deliberately perturbed. In both settings, we assess hallucination rates alongside overall response quality (details in §4.1).
@@
Based on our experiments, we find that (1) integrating KG information generally reduces factual errors, especially for queries requiring deeper reasoning; (2) relying solely on internal LLM reasoning strategies (e.g., Chain-of-Thought Kim et al. (2023) and Self-Consistency Wang et al. (2022)) can introduce biases and hallucinations; (3) subgraph-based methods often achieve the best performance for simpler query types; and (4) incorporating KGs effectively reduces hallucinations in LLMs even when the KG is partially contaminated.
@@ 2 OKGQA: An Open-ended Knowledge Graph Question-Answering Benchmark
OKGQA is a comprehensive benchmark designed to assess how effectively LLMs enhanced with KGs perform in open-ended, real-world-like question answering scenarios. Unlike existing benchmarks that focus primarily on closed-ended tasks, OKGQA presents diverse open-ended question types that mirror the variable nature of practical applications. As illustrated in Figure 1, given a complex query and its corresponding subgraph in a KG, the system must be capable of understanding the relationships within the data and performing human-like reasoning over the KG content to compose a paragraph-long answer. In the following section, we first describe our dataset construction, including query generation via LLM templates and KG subgraph extraction with PPR pruning. We then introduce OKGQA-P, a benchmark variant designed to evaluate model robustness under KG perturbations, detailing our perturbation methods and the metrics used to assess semantic and structural deviations. Due to page limitations, we provide additional extensions of our benchmark—including a multilingual setup and further analyses—in Appendix §B.
@@ 2.1 Dataset Construction
We utilize a template-based approach to generate a diverse set of queries using LLMs, including categories such as descriptive, explanatory, predictive, comparative, and critical queries. Specific templates and example queries are detailed in Table 1, with the corresponding prompts provided in Appendix D. To ensure that the generated queries reflect real-world complexity and relevance, we adopt an iterative optimization process that combines automated and human evaluations to refine the query generation process (see Appendix A.1 for details). Initially, we generate a diverse set of query candidates from a seed instruction. These candidates are then scored automatically by an LLM-based evaluator, which assigns quality scores $s_{\text{auto}}$ on a scale of 1-10, where higher scores indicate better performance across multiple criteria. Subsequently, human evaluators assess the same queries, producing normalized scores $s_{\text{human}}$ within the same range as $s_{\text{auto}}$. To refine the query generation, we iteratively optimize the input instructions by minimizing the discrepancy between $s_{\text{human}}$ and $s_{\text{auto}}$, thereby aligning automated assessments with human judgment. Additionally, the generated queries are categorized by complexity, with detailed statistics shown in Figure 2.
@@
To manage the size of KGs while covering relevant context for the queries, we follow previous work Yih et al. (2016); Talmor and Berant (2018) by sampling subgraphs from DBpedia (Noted that all queries in OKGQA can be answered using DBpedia). Specifically, we collect all triples within the $K$-hop neighborhood of entities mentioned in each query. We set $K=2$ to balance graph coverage and computational feasibility. As increasing beyond $2$-hop subgraphs generally leads to exponential growth in edges and nodes Jin et al. (2020), which increases excessive noise and complicates graph retrieval [^2] This choice aligns with common practices in benchmarks such as WebQSP Yih et al. (2016) and CWQ Talmor and Berant (2018), where 2-hop subgraphs are widely used for similar KGQA tasks.. To further reduce the size of subgraphs, we leverage Personalized Page-Rank (PPR) Bahmani et al. (2010) to prune the nodes/edges that are not relevant to the query (the details of the PPR algorithm are discussed in Appendix A.2). We compare the statistics of subgraphs before and after PPR pruning in Figure 2(a).
@@ 2.2 OKGQA-P: Benchmark with Noise & Perturbations in KGs
To quantify how much the perturbed KG has deviated from the original KG, i.e., $f(\mathcal{G},\mathcal{G}^{\prime})$, we leverage metrics from Raman et al. (2020) to evaluate semantics (ATS) and structural (SC2D, SD2) similarity between perturbed KG and original KG. Intuitively, ATS leverages a pre-trained LM for link prediction to measure the probability of each edge from $\mathcal{G}^{\prime}$ existing in $\mathcal{G}$, while SC2D and SD2 measure the structural similarity between two KGs based on local clustering coefficient and degree distribution. For each metric, higher value indicates higher similarity. The detailed description of these metrics can be found in Appendix A.5, with corresponding results shown in Figure 5.
@@ 3.1 Formalization
We formalize the KG-augmented framework as follows. Given a user query $q$, a pretrained language model generates a paragraph-like answer $a$ by modeling the conditional probability $p(a|q)$. To explore whether KGs help reduce hallucinations of LLMs, we introduce the retrieved knowledge $\mathcal{Z}$ from the KG and define:
@@ 3.2 Graph-guided retrieval (G-retrieval)
  • Sub-graph retrieval: building on He et al. (2024), we use the Prize-Collecting Steiner Tree (PCST) algorithm to find a connected subgraph $\mathcal{S}$ that maximizes $S(\mathcal{S})=\sum_{n\in V_{\mathcal{S}}}p_{v_{i}}+\sum_{e\in E_{\mathcal{S}}}p_{e_{i}}-\sum_{einE_{\mathcal{S}}}c_{e}$. Unlike in path-retrieval, we only yield one subgraph that maximizes the total score.
@@ 4 Experiments
In this section, we first introduce the evaluation metrics and experiment setup, and then focus on two main research questions and provide relevant analysis: RQ1: Can KGs reduce hallucination in LLMs? and RQ2: How are KG-Aware methods affected by noise/perturbations in KGs?
@@ 4.1 Evaluation Metrics & Setup
In addition to the hallucination metrics, we propose four metrics using LLM-as-evaluator (Li et al., 20 2) to quantify the quality of generated responses from LLM (Edge et al., 2024; Wang et al., 2023). In specific, we use G-Eval (Liu et al., 2023) framework for the evaluation and provide relevant Wikipedia pages of each query as context to enhance G-Eval’s robustness and stability. The four metrics are defined as follows: (1) Context Relevance: measures how well the generated response aligns with the provided context. (2) Comprehensiveness: assesses how thoroughly the answer addresses all aspects and details of the question. (3) Correctness: measures the clarity and specificity of the generated answer to the question. (4) Empowerment: evaluates how well the generated answer helps the reader understand the topic and make informed decisions. The corresponding prompts are detailed in Appendix D.
We use gpt-4o-mini (from November 2024 to January 2025) as LLM backbone for all the evaluation metrics. As using LLM-as-evaluator frameworks may raise concern regarding potential self-enhancement or bias from the selection of the backbone models Gu et al. (20 2); Li et al. (20 2), we conduct additional analysis in Appendix A.4 (including human evaluation alignment and cross-validation across different LLM backbones), and find that the choice of LLMs in the LLM-as-evaluator framework has little impact on the overall evaluation and the results demonstrate high correlation with the human evaluation, supporting the reliability of our testing.
For our experiments, we consider a range of widely used LLMs of different scales for testing, including GPT-4o, GPT-4o-mini (from November 2024 to January 2025), Llama-3.1-8B-instruct (Dubey et al., 2024), Mistral-7B-instruct-v0.3 (Jiang et al., 2023a), and Gemma-2-9B-it (Team et al., 2024). Considering the trade-off between cost and performance, we use text-embedding-3-small model from OpenAI (from November 2024 to January 2025) as embedding model for G-retrieval process. To ensure the reproducibility of the experiments, we set $\mathrm{temperature}=0.7$ and $\mathrm{top\_p}=1.0$ for all models. We use the API service from OpenAI [^3] https://openai.com/ and OpenRouter [^4] https://openrouter.ai/ for our experiments which host detailed snapshots of the utilized model versions.
@@ 4.2 RQ1: Main Results - Can KGs Reduce Hallucination in LLMs?
Directly performing reasoning in the LLM itself does not mitigate hallucinations. We benchmark the hallucination ratio of LLMs using internal reasoning strategies like CoT and Self-consistency. As shown in Var-1 in Table 2, these methods can improve response quality (i.e., G-Eval) compared to baselines, but do not consistently improve factuality, and sometimes even diminish. This shows that relying solely on internal reasoning is inadequate for mitigating hallucinations, highlighting the necessity for external knowledge to address this issue effectively.
@@ 4.3 RQ2: How Are KG-Aware Methods Affected by Noise / Perturbations in KGs?
We test different KG-augmented LLMs on our OKGQA-P setting, where we deliberately perturb and contaminate the semantics and structure of KGs to simulate the real-world situation where KGs may not have high quality. Specifically, we consider different perturbation methods discussed in §2.2 and control the perturbation level based on the percentage of KG edges being perturbed. We first illustrate how much the perturbed KG has been deviated from the original KG with the increase of perturbation level, shown in Figure 5. It shows that the perturbation methods like edge deletion, rewiring and swapping have relatively weak influence on ATS (which intuitively measures semantic similarity), even as the perturbation level increases. For the edge deletion methods, only if the perturbation level reaches 1.0, the ATS goes to 0, otherwise, the ATS remains higher compared to other settings.
Figure 6 illustrates the hallucination ratio using different methods on the perturbed KGs. We observe that (1) FS-SG consistently outperforms FS-T and FS-P even at higher perturbation levels, demonstrating its robustness by maintaining higher scores as perturbations increase. (2) FS-T and FS-P exhibit similar trends, each showing a significant performance drop as perturbation levels increase. Particularly, performance of FS-T and FS-P deteriorate when the perturbation level reaches 50%, i.e., becoming worse than the baseline using CoT. (3) On the setting using Relation Replacement which severely harms the semantics of the KGs, FS-T and FS-P decline more sharply than FS-SG. However, it still outperforms the baseline when the perturbation level is smaller than 40%.
In summary, we find that the effectiveness of KG-derived information diminishes with a perturbation level at 50%, surpassing this level leads to a further decrease in performance. We think that before this perturbation level at 50%, incorporating external knowledge from KGs can mitigate hallucinations in LLMs compared to baseline using CoT. Considering practical scenarios, platforms like Wikidata are less likely to have perturbations as severe as 50% due to their ongoing updates and community-based quality control. This ensures the relevance and applicability of our findings in real-world settings.